# Limitations of Using Identical Distributions for Training and Testing When Learning Boolean Functions

### v1

@@ Abstract
When the distributions of the training and test data do not coincide, the problem of understanding generalization becomes considerably more complex, prompting a variety of questions. In this work, we focus on a fundamental one: Is it always optimal for the training distribution to be identical to the test distribution? Surprisingly, assuming the existence of one-way functions, we find that the answer is no. That is, matching distributions is not always the best scenario, which contrasts with the behavior of most learning methods. Nonetheless, we also show that when certain regularities are imposed on the target functions, the standard conclusion is recovered in the case of the uniform distribution.
@@ 1 Introduction
In this work, however, we do not attempt to establish under which shifts certain learning methods are robust, but rather address a more fundamental question: is it really optimal for the test distribution to be identical to the training distribution? To investigate this, we use tools from computational learning theory. In particular, instead of adopting the traditional PAC-learning framework introduced by Valiant [valiant1984theory], we focus on a variant in which the efficient evaluation of the model is also considered [gyurik2023exponential, perez2024classical]. Intuitively, one might think the answer to the question is trivially “yes.” However, we show that this is not generally the case, assuming that one-way functions exist, which highlights the extremely complex and counterintuitive nature of understanding generalization in this setting. Nonetheless, assuming certain regularities, we can guarantee that this effect disappears for the uniform distribution.
@@ 5 Conclusions
In summary, we have shown that, under the assumption that one-way functions exist, having the training distribution equal to the test distribution is not always optimal. Furthermore, we have shown that this effect appears even when efficiently samplable sequences of distributions are considered. However, to achieve this latter result, the constructed example relies on the fact that some pairs $(x,c_{n}(x))$ are much more informative than others. To study the scenario in which this is not the case, we introduce the concept of a regular concept class, in which all pairs are similarly informative. For this type of concept class, we conclude that, at least for the uniform distribution, the counterintuitive results of Theorem 1 and Theorem 2 do not hold. Finally, as future work, it would be interesting to study other types of concept classes that exhibit the same property as regular concept classes but for a wider set of distributions.

### v2

@@ Abstract
When the distributions of the training and test data do not coincide, the problem of understanding generalization becomes considerably more complex, prompting a variety of questions. Prior work has shown that, for some fixed learning methods, there are scenarios where training on a distribution different from the test distribution improves generalization. However, these results do not account for the possibility of choosing, for each training distribution, the optimal learning algorithm, leaving open whether the observed benefits stem from the mismatch itself or from suboptimality of the learner. In this work, we address this question in full generality. That is, we study whether it is always optimal for the training distribution to be identical to the test distribution when the learner is allowed to be optimally adapted to the training distribution. Surprisingly, assuming the existence of one-way functions, we find that the answer is no. That is, matching distributions is not always the best scenario. Nonetheless, we also show that when certain regularities are imposed on the target functions, the standard conclusion is recovered in the case of the uniform distribution.
@@ 1 Introduction
Interestingly, as pointed out in [gonzalez2015mismatched, canatar2021out], shifts in the distributions are not always detrimental. In particular, these works show that for certain fixed learning methods, there exist scenarios where a mismatch between the training and test distributions can actually improve generalization. However, they do not consider selecting the optimal learning method for each training distribution. Thus, the observed advantages of mismatched distributions in those settings may simply arise from the use of suboptimal learners rather than from the mismatch itself.
In this work, we address the fundamental question: is it really optimal for the test distribution to be identical to the training distribution when the learner is allowed to be optimally adapted to the training distribution? To investigate this, we use tools from computational learning theory. In particular, instead of adopting the traditional PAC-learning framework introduced by Valiant [valiant1984theory], we focus on a variant in which the efficient evaluation of the model is also considered [gyurik2023exponential, perez2024classical]. Intuitively, one might think the answer to the question is trivially “yes.” However, we show that this is not generally the case, assuming that one-way functions exist, which highlights the extremely complex and counterintuitive nature of understanding generalization in this setting. Nonetheless, assuming certain regularities, we can guarantee that this effect disappears for the uniform distribution.
@@ 5 Conclusions
In summary, we have shown that, under the assumption that one-way functions exist, having the training distribution equal to the test distribution is not always optimal, even when the learner is allowed to be optimally adapted to the training distribution. Furthermore, we have shown that this effect appears even when efficiently samplable sequences of distributions are considered. However, to achieve this latter result, the constructed example relies on the fact that some pairs $(x,c_{n}(x))$ are much more informative than others. To study the scenario in which this is not the case, we introduce the concept of a regular concept class, in which all pairs are similarly informative. For this type of concept class, we conclude that, at least for the uniform distribution, the counterintuitive results of Theorem 1 and Theorem 2 do not hold. Finally, as future work, it would be interesting to study other types of concept classes that exhibit the same property as regular concept classes but for a wider set of distributions.